OpenStudy (bj11684):

After a concert for 1,200 people, only 9 people said they would not go to see the band again. What percent of the people who went to the concert said they would not go to see the band again?

Atsie (atsie):

Whats your guess?

OpenStudy (calculusxy):

\(\text{___% of 1,200 = 9}\) since we don't know the percentage, i just left it blank. do you understand so far?

OpenStudy (bj11684):

yes

OpenStudy (calculusxy):

so it's basically like: \(\large \color{red}{\text{Percent}} \text{ } \text{of} \text{ } \color{blue}{\text{Whole}} = \color{green}{\text{Part}}\)

OpenStudy (anonymous):

Hey calculusxy. Were you studying for the Hunter Entrance Exam

OpenStudy (calculusxy):

Now we can set up the proportion to solve for the percent: \(\large \frac{Part}{Whole} = \frac{Percent}{100}\) since we don't know the percent, it will be left as \(x\) we can fill this proportion up by filling what we know \(\large \frac{9}{1,200} = \frac{x}{100}\) you can solve for \(x\) in this proportion

OpenStudy (anonymous):

calculusxy are you listening

OpenStudy (calculusxy):

Yes i did @infinityknowledge

OpenStudy (bj11684):

ok

OpenStudy (anonymous):

Were you studying for the Hunter Entrance Exam

OpenStudy (calculusxy):

Yes i did study for the hunter entrance exam two years back @infinityknowledge

OpenStudy (anonymous):

Can you help. Im preparing for it. I saw your questions

OpenStudy (calculusxy):

@bj11684 What did you get for x?

OpenStudy (bj11684):

wait

OpenStudy (bj11684):

x=3/4

563blackghost (563blackghost):

close but incorrect...if it was 3/4 than 75% would have not attended it again and according to your question only 9 people said no so it would be alot less than 75%...

OpenStudy (bj11684):

it is 0.75%

563blackghost (563blackghost):

OMG I didnt see your answer right...You are actually correct ^^ sorry...

563blackghost (563blackghost):

It is 0.75%....great job ^^

563blackghost (563blackghost):

Please close your question once you get the help you needed....